verify that -2x^2y+y^2=1 is the implicit solution of the of the differential equation (x^2-y)dy/dx + 2xy=0


1
Expert's answer
2022-02-14T17:31:20-0500

Let us verify that "-2x^2y+y^2=1" is the implicit solution of the of the differential equation "(x^2-y)\frac{dy}{dx} + 2xy=0." For this let us find the implicit derivative of "-2x^2y+y^2=1." It follows that

"-4xy-2x^2\frac{dy}{dx}+2y\frac{dy}{dx}=0," that is equivalent to "2xy+x^2\frac{dy}{dx}-y\frac{dy}{dx}=0." The last equation is equivalent to "(x^2-y)\frac{dy}{dx} + 2xy=0." Consequently, "-2x^2y+y^2=1" is indeed the implicit solution of the of the differential equation "(x^2-y)\frac{dy}{dx} + 2xy=0."